AnimeGhoul8863:

After a dreary day of rain, the sun peeks through the clouds and a rainbow forms. You notice the rainbow is the shape of a parabola. The equation for this parabola is y = -x2 + 36.

AnimeGhoul8863:

1 attachment
AnimeGhoul8863:

In the distance, an airplane is taking off. As it ascends during take-off, it makes a slanted line that cuts through the rainbow at two points. Create a table of at least four values for the function that includes two points of intersection between the airplane and the rainbow. Analyze the two functions. Answer the following reflection questions in complete sentences. What is the domain and range of the rainbow? Explain what the domain and range represent. Do all of the values make sense in this situation? Why or why not? What are the x- and y-intercepts of the rainbow? Explain what each intercept represents. Is the linear function you created with your table positive or negative? Explain. What are the solutions or solution to the system of equations created? Explain what it or they represent. Create your own piecewise function with at least two functions. Explain, using complete sentences, the steps for graphing the function. Graph the function by hand or using a graphing software of your choice (remember to submit the graph).

AnimeGhoul8863:

@dude e,e can i getz some help preaze

woolyfrog:

lol look at my question

AnimeGhoul8863:

In the distance, an airplane is taking off. As it ascends during take-off, it makes a slanted line that cuts through the rainbow at two points. Create a table of at least four values for the function that includes two points of intersection between the airplane and the rainbow. -6 6 0 0

AnimeGhoul8863:

(-6,0) (6,0)

dude:

Yeah 2 posts have already been made for this question idk where you guys would like me to answer

dude:

This include much more information though Anyway you dont need the zeroes of the rainbow for the table

AnimeGhoul8863:

if you dont have to add the 0 then i have to find 2 more numbers e.e hmmmm

woolyfrog:

so you chose the points -6,0 to 6,0

AnimeGhoul8863:

yes cause it says it makes a slanted line that cuts through the rainbow at two points and if you look at the graph

AnimeGhoul8863:

it has 2 points -6,0 and 6,0

dude:

Well yes but it’s an airplane, if we use those points the plane would be on the ground the whole time

woolyfrog:

can I use -6,0 to 0,36

AnimeGhoul8863:

would (-6,0) and (11,4) work?

dude:

-6,0 to 0,36 is okay 11, 4 is not on the graph but it works anyway

dude:

Anyway to create the equation of the line, use the slope between two points formula \(\large \frac{\Delta y}{\Delta x}\) or \(\large \frac{y_2-y_1}{x_2-x_1}\)

dude:

Or rise/run

AnimeGhoul8863:

i thought of it in the way planes dont go straight up when they lift off they go diagonal then up so 11,4 would make it look like the airplane is rising

dude:

Yeah it works, I mean if you wanted to find the slope visually it might be harder

AnimeGhoul8863:

so it would be better and easier to use the 0,36?

dude:

I would suggest that you both choose different points (Teachers might catch on similar points)

woolyfrog:

@dude

woolyfrog:

I'm confused on what you said about creating the equation of the line

woolyfrog:

@Marziman

woolyfrog:

please help

AnimeGhoul8863:

im so sorry dude my laptop screen went black and i had to restart it e.e

dude:

@woolyfrog Do \(\large \frac{36-0}{0--6}\)

dude:

Its okay dont worry

woolyfrog:

ok thank you

AnimeGhoul8863:

so mine would look like \[\frac{ 11-4 }{ 0- -6 }\]

dude:

Close \(\large \frac{4-0}{11--6}\)

woolyfrog:

ok I wrote it as 36-0/0--6 what is next

AnimeGhoul8863:

so the answer to this question...in the distance, an airplane is taking off. As it ascends during take-off, it makes a slanted line that cuts through the rainbow at two points. Create a table of at least four values for the function that includes two points of intersection between the airplane and the rainbow. would be \[\frac{ 4-0 }{ 11--6 }\]

AnimeGhoul8863:

wait no it wouldnt cause they are asking for a table

AnimeGhoul8863:

x | y --|---- 4 | 11 0 | -6

AnimeGhoul8863:

like this? im guessing

dude:

We're finding the equation of the line so we can just substitute any number and use it on the table But yes that table looks good

dude:

Solve it wooly

AnimeGhoul8863:

so would this table be the answer to the first question?

woolyfrog:

x | y --|---- 36 | 0 0 | 6

woolyfrog:

@dude

AnimeGhoul8863:

wooly dont keep tagging him he will be back you just have to wait tagging him constantly will only frustrate and annoy the person helping e.e we gots dis so dont stress

woolyfrog:

Hai okay

dude:

Oh whoops I have to correct your tables x | y -- |---- -6 | 0 11 | 4 and x | y --|---- -6 | 0 0 | 36 The directions say you need 4 points, these are only 2 Find the slopes in each

AnimeGhoul8863:

Create a table of at least four values for the function that includes two points of intersection between the airplane and the rainbow.

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @woolyfrog can I use -6,0 to 0,36 \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

it calls for 2 points and both points contain 2 values i think thats what i means e.e

AnimeGhoul8863:

but i could be wrong x.x

dude:

Well I guess that works too but your point 11,4 isn't intersecting the graph (I think it would be best to include at least 4 points just to be sure)

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude @woolyfrog Do \(\large \frac{36-0}{0--6}\) \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

............. T~T fineeeee but how do we get 4 point the airplane can only go in one direction

woolyfrog:

how do I found 2 more values

dude:

Well first solve that so we can just get values easier

dude:

Anime \(\large \frac{4-0}{11--6}\)

woolyfrog:

36/6

dude:

36/6 is right, you can simplify that further

woolyfrog:

4/17

dude:

No

dude:

\(\frac{36}{6}=6\)

woolyfrog:

Hai yes

AnimeGhoul8863:

4-0/11--6= 4/17?

dude:

Oh that is Anime's

dude:

Yeah

dude:

Wooly your equation is \(y=6x+b\) Anime your equation is \(y=\frac{4}{17}x+b\) Both of you substitute -6 for x and 0 for y

dude:

And solve for b

woolyfrog:

0 = -36 + b

dude:

Yes, and what is b?

woolyfrog:

36

AnimeGhoul8863:

b=24/17?

dude:

Yep wooly = y=6x+36 Anime = y = \(\frac{4}{17}x+\frac{24}{17}\) Now to find your other points, choose any number and substitute it into x

woolyfrog:

any number?

AnimeGhoul8863:

would 3 work as x e.e

dude:

Yeah any number 3 is good

dude:

Uh I okay let me clarify Any point between -6 and 6

dude:

wooly you cannot use -6 and 0 (because you already know those points)

dude:

wooly you can use -5,-4,-3,-2,-1, 1, 2,3,4,5,6

AnimeGhoul8863:

do we simplify?

dude:

Anime you can use -5,-4,-3,-2,-1,0,1,2,3,4,5,6

dude:

Yep

AnimeGhoul8863:

y=36/17

woolyfrog:

y=6 2 +36 y= 48

dude:

Yep Anime and now do one more number

dude:

Which number did you do wooly? The math looks wrong if you did x=2

AnimeGhoul8863:

like add another number like instead of 3 would 7 work?

woolyfrog:

I did 2

AnimeGhoul8863:

using 3 = 36/17 Using 7= 52/17

dude:

Actually Anime do 5.765 (They are asking for 2 points that intersect) Are you allowed to use decimals?

AnimeGhoul8863:

ummmm idk tbh

dude:

Okay wooly do another point

AnimeGhoul8863:

i mean i can try e.e

dude:

You chose slightly hard points to work witb

dude:

Uh I'll give you an easier table anime

woolyfrog:

y=6 3 +36 y= 54

AnimeGhoul8863:

alright e.e

dude:

Wooly your table is x | y --|---- -6 | 0 0 | 36 2 | 48 3 | 54

dude:

wooly, make a separate post so I can help you with your domain and range if you need

dude:

Anime we can use points x | y -- |---- -3 | 27 2 | 32 Equation of this line is \(y=x+30\) Now substitute 2 x values into this

AnimeGhoul8863:

sorry i got confused ....hold on

dude:

Okay so I completely changed your table (because I assume you do not like decimals and fractions)

AnimeGhoul8863:

yeah i kinda have a thing against fractions XD but so this whole thing is gone x | y --|---- 4 | 11 0 | -6

AnimeGhoul8863:

are we starting from scratch now?

AnimeGhoul8863:

or are we keeping these and changing the 2 other points

dude:

Yeah just forget those points x | y --|---- 4 | 11 0 | -6 This is from scratch

dude:

x | y -- |---- -3 | 27 2 | 32 Equation of this line is \(y=x+30\) Use this one (it is much easier to work with)

AnimeGhoul8863:

im sorry dude im still confused hold on where are these points on the graph?

dude:

1549301005-5c58750d78b49562bba49000-image.png

AnimeGhoul8863:

ok

dude:

So the line is 1549301113-5c587579695aac2b41e5b800-image.png

dude:

Okay so now substitute any 2 numbers into x (I recommend x=0 and x=-1)

dude:

Which would be \(y=(0)+30\) and \(y=(-1)+30\)

AnimeGhoul8863:

are we finding the y intercept?

dude:

Its the easiest point to find so yes

AnimeGhoul8863:

y=0+30= b=30 y=-1+30=b=29

dude:

Yep so your table is x | y -- |---- -3 | 27 -1 | 29 0 | 30 2 | 32

AnimeGhoul8863:

alright what is next then

dude:

`What is the domain and range of the rainbow? Explain what the domain and range` `represent. Do all of the values make sense in this situation? Why or why not?` `What are the x- and y-intercepts of the rainbow? Explain what each intercept represents.` Make a new post for this, its already long enough

AnimeGhoul8863:

okie